# Self-Calibrating Gaussian Splatting for Large Field of View Reconstruction

### v1

@@ Abstract
In this paper, we present a self-calibrating framework that jointly optimizes camera parameters, lens distortion and 3D Gaussian representations, enabling accurate and efficient scene reconstruction. In particular, our technique enables high-quality scene reconstruction from Large field-of-view (FOV) imagery taken with wide-angle lenses, allowing the scene to be modeled from a smaller number of images. Our approach introduces a novel method for modeling complex lens distortions using a hybrid network that combines invertible residual networks with explicit grids. This design effectively regularizes the optimization process, achieving greater accuracy than conventional camera models. Additionally, we propose a cubemap-based resampling strategy to support large FOV images without sacrificing resolution or introducing distortion artifacts. Our method is compatible with the fast rasterization of Gaussian Splatting, adaptable to a wide variety of camera lens distortion, and demonstrates state-of-the-art performance on both synthetic and real-world datasets. More details and videos can be found at the project page: https://denghilbert.github.io/self-cali/.
@@ 1 Introduction
Recent advances in 3D scene reconstruction, such as Neural Radiance Fields (NeRFs) [48] and Gaussian Splatting (3DGS) [37], have shown remarkable success for high quality novel-view synthesis (NVS). However, these methods typically require dense image captures, relatively narrow FOV, and precise camera pose estimation from structure-from-motion (SfM) techniques [58, 59]. In practical applications, large FOV lenses such as fisheye lenses are often used in fields like robotics and virtual reality [18, 40] because they capture larger regions of a scene with fewer images, enabling faster data collection and reconstruction. Yet, these lenses introduce significant geometric distortion, making accurate camera calibration and scene reconstruction challenging.
Sometimes, the intrinsic properties of wide-angle imaging systems can be pre-calibrated using specialized setups such as a room covered in calibration markers. But these processes can be slow and cumbersome. Polynomial camera distortion models, such as those implemented in OpenCV [10, 58], work well for standard lenses but struggle to modeling the complex and non-linear distortions of fisheye optics. As a result, many frameworks require fisheye images to be “undistorted” into perspective images before reconstruction, resulting in images with significant stretching and/or cropping relative as seen in Fig. 2.
To address these challenges, we explore self-calibration techniques that jointly optimize camera parameters with 3D scene representations. Such approaches have roots in methods that calibrate cameras without explicit calibration targets [67, 53, 15, 16]. More recently, self-calibration has been integrated into radiance fields and 3D Gaussian frameworks [49, 34, 46], where the photometric loss is used to refine both extrinsic and intrinsic parameters. However, these methods still lack the ability to accurately model the complex lens distortions typical for wide FOV imagery. This results in significant misalignment artifacts, especially in the peripheral regions.
In this work, we introduce Self-Calibrating 3D Gaussian Splatting, a differentiable rasterization pipeline that optimizes both camera lens distortion and scene representations using 3D Gaussians. Our approach is designed to be expressive enough to model diverse distortions while remaining computationally efficient and well regularized for stable training. Previous methods have attempted to handle lens distortion using grid-based ray offsets [34] or invertible neural networks [62]. While grid-based models are fast, they often produce noisy and unstable results when supervised solely by photometric loss. In contrast, neural network–based approaches offer better regularization but are prohibitively expensive when applied to each individual Gaussian.
@@
To overcome these limitations, we propose a novel hybrid neural field that combines the strengths of both methods, striking a balance between expressiveness and computational efficiency, as shown in Fig 3. Specifically, our method adopts invertible Residual Networks [8] to predict displacement on a normalized sparse grid, followed by bilinear interpolation to generate a continuous distortion field.
For images with large fields of view, using a naive single planar projection results in severe pixel stretching and distortion, particularly in the peripheral regions. To mitigate this, we introduce a novel resampling strategy based on a cubemap representation, which evenly distributes image data across cubemap faces. This approach significantly reduces distortion artifacts and provides a clear, relatively undistorted omnidirectional view of the scene, preserving pixel resolution even in the outer areas.
@@
To validate the efficacy of our method, we conduct extensive experiments on both synthetic datasets and several real-world scenes, including the FisheyeNeRF datasets [34] and our synectics Mitsuba dataset. Our system effectively calibrates camera parameters and lens distortion, achieving superior Gaussian Splatting performance compared to existing methods when using uncalibrated fisheye cameras. Importantly, our parametrization is not limited to a single fisheye camera model; rather, it is designed to be flexible and adaptable, accommodating a wide variety of camera models and real-world distortions seamlessly, all without the need for pre-calibration. This flexibility enables our method to fully leverage the unique capabilities of each available lens, enabling rapid scene capture with comprehensive coverage.
@@ 2 Related Work
Lens distortion is an inherent property of all cameras. In general, non-linear distortion can be formulated as:
where K and are the intrinsic and extrinsic, respectively. denotes the pinhole projection, including dehomogenization to obtain 2D points on the image plane. The distortion model is parameterized as a polynomial function of the radial distance:
where , , , are the parameters of the Brown–Conrady model [17, 11], derived from calibration, and . Scaramuzza et al. [57] first proposed a unified model for larger FOV fisheye lenses, which has been adopted in several works [12, 42]. The most widely used fisheye model [32] describes the distortion as a function of the angular distance from the projection center:
where , and the distances of projected points to the image plane are normalized to 1. The 3D Gaussian [37] assumes a standard perfect pinhole camera model and typically requires COLMAP [58] to undistort images before reconstruction. To relax this constraint, recent methods [46, 49] adopt parametric models like Eq. 3 to extend 3D Gaussian Splatting techniques to fisheye images. However, these methods still rely heavily on camera calibration for accurate estimation and fix the projection in rasterization, making it difficult to generalize to various types of cameras. Works like [49] introduce ray tracing into the rasterization pipeline and approximate Gaussians bounding with icosahedron which potentially jeopardize the efficiency of rasterization. Some works [45, 4] also explore reconstruction from omnidirectional 360 panoramas, but the major difference is that panoramas require calibration to stitch two fisheye images and do not preserve raw geometric consistency at the stitching boundary. We address the aforementioned limitations by applying a hybrid distortion field that is compatible with the 3D Gaussian Splatting pipeline. Our experiments demonstrate existing methods such as Fisheye-GS [46] and ADOP [55] adopt traditional camera distortion models, such as those in Eqs. 2 and 3 into the rasterization process, are not expressive enough to handle the distortions present in large FOV cameras.
@@
The bundle adjustment process can be extended to include optimization of camera lens parameters as well as the poses. This way cameras do not need to be calibrated in a separate operation; this process is known as self-calibration [68, 54, 19, 30]. Camera calibration without a known calibration target is challenging as it relies on strong assumptions about the scene structures and geometric priors for establishing reliable 3D-2D correspondence [6, 14, 2]. Camera auto-calibration methods [68, 54, 19] extend this idea to derive camera intrinsics from multi-view observations of unstructured scenes, which was further advanced in recent studies [22, 21, 28]. Typically, the camera calibration process involves careful selection of appropriate camera lens models from a large family of high-order polynomial models and camera projection functions [23, 56, 36, 20, 60]. On the other hand, non-parametric models have been developed for broad applicability across different camera and lens combinations [25, 13, 29, 44], while additional regularization was needed to ensure the underlying distortion to be smooth [51]. With advances in differentiable rendering and rasterization pipelines [41, 65], recent works [55, 62, 34] showed that camera lens distortion can be optimized together with other parameters through a differentiable projection module. Prior works also adapt NeRF for panoramic inputs and fisheye distortions [27, 31, 63, 43]. These solutions usually use parametric models tailored for specific lenses of interest and are not generalizable to a wide variety of lenses types. SCNeRF [34] proposes to model a residual projection matrix and residual raxel parameters [26], which are interpolated on a sub-sampled pixel grid. Cam-P [52] proposes a novel preconditioner for the pose optimization, further improving the modeling of camera extrinsics and intrinicics jointly with NeRF. NeuroLens [62] optimizes lens parameters through an invertible neural network. This work builds on the insight of prior self-calibration methods and introduces a novel and efficient way of modeling lens distortion integrated with Gaussian Splatting.
@@ 3 Method
Given uncalibrated wide-angle captures, we aim to develop an algorithm that produces high-quality reconstructions using 3D Gaussians. Our method is designed to be robust to severe distortion in peripheral regions of images and various wide-angle lens effects. We achieve this through two key steps. First, we extend Gaussian splatting to support a broader range of camera models, such as fisheye lenses, as discussed in Sec. 3.2. To model lens distortion, we apply a hybrid distortion field to diverse real-world scenarios where cameras with various distortions are used. Second, we replace the traditional single planar projection in 3DGS with cubemaps and introduce the depth sorting strategy accordingly, as detailed in Sec. 3.3.
@@ 3.1 Background of Gaussian Splatting
3DGS explicitly represents a 3D scene with a collection of Gaussians, each parameterized by a center , a scaling vector , and a quaternion representing the orientation. Each Gaussian can be formulated as:
where is a diagonal matrix with the vector on the diagonal, and is the rotation matrix for the quaternion . Each Gaussian is also associated with a -dimensional vector that stores spherical harmonic coefficients to represent the color of the Gaussian.
To produce an image from this set of Gaussians, we need to assume a projection function that takes a 3D point and projects it to pixel coordinates conditioned on camera parameters . Under this projection function, the 3D Gaussians are projected to 2D for rasterization [69]:
where is the affine approximation of the projection at point while parameterized by . In the original 3DGS paper, the authors assume perspective pinhole camera models free of lens distortion. The final image for a set of Gaussians is produced using alpha blending [41, 65]. Specifically, the RGB color for pixel location is given by the following equation:
where is an ordered set of indices for relevant Gaussians to pixel location sorted by depth, indicates the first element of , is the view-dependent RGB color, is the opacity given by applying Eq. 5 on the 2D Gaussians. Finally, this set of 3DGS is optimized using a photometric loss:
where is the loss weight, is the L1 loss and is the D-SSIM [5] loss. The parameters of Gaussians are usually optimized using gradient-based algorithms together with adaptive control. [37].
@@ 3.2 Lens Distortion Modeling
In this section, we extend the Gaussian splatting technique to accommodate a broader class of camera lenses, including fisheye and wide-angle cameras by modeling the lens distortion. Lens distortion is typically captured by a distortion function defined in camera coordinates. A Distortion function parameterized by maps pixel locations from a rectified image to locations of a distorted image. Ideally, the mapping should be: 1) expressive enough to model various lens distortions, 2) well regularized such that it can be optimized together with the 3D scenes, and 3) efficient so that it does not add significant overhead. While existing methods have explored using parametric camera models, grid-based methods, and deep-learning methods, none of these methods perfectly satisfy all three criteria.
@@
where the optimizable parameters are a grid storing 2D vectors representing the distortion. Bilinear interpolation function , with being the bilinear interpolation weights of location . Such a grid-based method is expressive and efficient, since is sparse (i.e., weights are zeros except for the those in the neighbor nodes of ) and one can increase the grid resolution to model more complex functions. The grid-based method, however, lacks the proper smoothness needed to model lens distortion, resulting in overfitting and suboptimal solutions (Fig. 3 Top).
@@
An alternative way to model the distortion is using a neural network with appropriate inductive bias. NeuroLens [62] proposes to use invertible residual networks [8] to represent non-linear lens distortions as a diffeomorphism. Specifically, the deformation mapping is modeled by a residual network:
where is a neural network parameterized by with Lipschitz constant bounded by (i.e., for all , , and ). represents a residual block with 4 linear layers. L denotes the number of blocks, which is 5 in our case, and the circle denotes function composition. Such constraints make the network invertible, and its inverse can be obtained using a fixed-point algorithm [8].
While an invertible ResNet provides both expressivity (i.e., being able to fit many lenses) and regularity, it is computationally expensive to deploy it directly to 3DGS for several reasons. To compute the alpha-blending weights and when rendering an image in Eq. 6, forward and backward passes of for each Gaussian are required. This is computationally expensive for large-scale scenes with millions of 3D Gaussians, and motivates us to find a more efficient solution that can leverage the inherent inductive bias of the invertible ResNet. We show more ablation comparisons in the supplementary to demonstrate the effective regularization of iResNet compared with regular ResNet.
@@
Given that the grid-based method is efficient yet tends to overfit, while the invertible ResNet has an appropriate inductive bias but is not efficient, we propose a hybrid method that combines the advantages of both. Specifically, we use the invertible ResNet to predict the flow field on a sparse grid and apply bilinear interpolation:
where is a sparse grid of fixed control points (pixel locations and is the resolution of control points instead of images) and is an invertible ResNet parameterized by .
Unlike existing hybrid neural fields [50] where networks are applied after grid interpolation, our approach uses invertible Residual Networks to predict displacement vectors on a sparse grid where bilinear interpolation is applied to produce a continuous displacement field, shown at the bottom of Fig. 3. The advantage of this architecture is that we only need to compute the expensive forward and backward ResNet mappings for locations at . This scales with the grid resolution, independent of the number of Gaussians in the scene. The additional operation that needs to be done for each Gaussian is , which is affordable and parallelizable.

### v2

# Self-Calibrating Gaussian Splatting for Large Field-of-View Reconstruction

@@ Abstract
Large field-of-view (FOV) cameras can simplify and accelerate scene capture because they provide complete coverage with fewer views. However, existing reconstruction pipelines fail to take full advantage of large-FOV input data because they convert input views to perspective images, resulting in stretching that prevents the use of the full image. Additionally, they calibrate lenses using models that do not accurately fit real fisheye lenses in the periphery. We present a new reconstruction pipeline based on Gaussian Splatting that uses a flexible lens model and supports fields of view approaching 180 degrees. We represent lens distortion with a hybrid neural field based on an Invertible ResNet and use a cubemap to render wide-FOV images while retaining the efficiency of the Gaussian Splatting pipeline. Our system jointly optimizes lens distortion, camera intrinsics, camera poses, and scene representations using a loss measured directly against the original input pixels. We present extensive experiments on both synthetic and real-world scenes, demonstrating that our model accurately fits real-world fisheye lenses and that our end-to-end self-calibration approach provides higher-quality reconstructions than existing methods. More details and videos can be found at the project page: https://denghilbert.github.io/self-cali/.
@@ 1 Introduction
Large field-of-view (FOV) lenses, such as fisheye lenses, are widely used in robotics [20], virtual reality [41], and autonomous driving [19] because they capture scenes with fewer images, enabling efficient data acquisition for reconstruction and novel view synthesis (NVS)[48]. However, most modern 3D reconstruction systems based on Neural Radiance Fields (NeRFs)[50] and Gaussian Splatting (3DGS) [38] cannot be directly applied to this type of input because they rely on perspective projection. As a result, fisheye images are typically re-projected into linear perspective images, and the resulting non-uniform sampling and stretching pose challenges for accurate camera calibration and scene reconstruction from wide-angle imagery.
Traditionally, imaging systems are pre-calibrated using specialized setups, such as calibration checkerboards. During calibration, polynomial distortion models [10, 59] are estimated and then used to re-project raw fisheye images into perspective ones for scene reconstruction. There are three important problems with this practice. First, resampling wide-field images leads to severe stretching and non-uniform sampling, as shown in Fig. 2(a), where perspective images exhibit far less uniform sampling of directions than fisheyes. Second, traditional parametric distortion models lack sufficient expressiveness to accurately model large FOV (e.g., 180) lenses. Third, pre-calibration prevents lens parameters from being optimized end-to-end along with reconstruction, meaning the final reconstruction does not fully minimize the reconstruction loss.
To avoid the separate pre-calibration stage, many recent works [51, 36, 47] integrate distortion modeling [66, 55, 15, 16] and optimization into NeRF and 3DGS frameworks. However, even when using photometric loss to refine distortion parameters alongside reconstruction, these approaches still exhibit significant misalignment in peripheral regions due to their reliance on traditional distortion models and single-plane projection. Both issues constrain the ability of current reconstruction frameworks to represent wide-FOV optical systems.
In this work, we introduce Self-Calibrating Gaussian Splatting, a differentiable rasterization pipeline that jointly optimizes lens distortion, camera intrinsics, camera poses, and scene representations using 3D Gaussians. Our approach effectively addresses the three aforementioned challenges, achieving high-quality reconstruction from large-FOV images without requiring resampling, pre-calibration, or polynomial distortion models.
@@
To improve lens modeling, we replace the conventional parametric distortion model with a novel hybrid neural field that balances expressiveness and computational efficiency, as illustrated in Fig. 3. Our method uses invertible residual networks [8] to predict displacements on a normalized sparse grid, followed by bilinear interpolation to generate a continuous distortion field.
To mitigate the stretching caused by resampling to a single-plane perspective, we use cubemap sampling as in [17] (Fig. 2(b)), which significantly reduces stretching and yields a more uniform pixel density, even in peripheral areas. Gaussians project into each limited-FOV face of the cubemap with bounded distortion.
Finally, our method combines lens calibration with bundle adjustment and scene reconstruction into an integrated end-to-end self-calibration process that minimizes the rendering loss over all unknowns jointly, leading to lower reconstruction errors compared to pre-calibration of distortion and/or camera pose.
@@
To validate our method, we conduct extensive experiments on both synthetic datasets and real-world scenes, including the FisheyeNeRF datasets [36], our own real-world captured scenes, and a synthetic dataset. Our system effectively calibrates extrinsics, intrinsics, and lens distortion, achieving better reconstruction performance compared to existing methods using uncalibrated fisheye cameras. Importantly, our system is not constrained to a single fisheye camera model; rather, it is designed to be flexible and adaptable, accommodating a wide range of cameras, from perspective to extreme fisheye, without requiring pre-calibration. This flexibility enables our method to fully leverage the unique capabilities of available lenses, ensuring comprehensive scene coverage and high-quality reconstructions.
@@ 2 Related Work
Lens distortion is an inherent property of all cameras. In general, nonlinear distortion can be formulated as:
where K and represent the intrinsic and extrinsic parameters, respectively. X is a 3D point in world coordinates. denotes the pinhole projection, including dehomogenization to obtain 2D points on the image plane. The distortion model is parameterized as a polynomial function of the radial distance:
where , , , are the parameters of the Brown–Conrady model [18, 11], derived from calibration, and . Scaramuzza et al.[58] first proposed a unified model for large-FOV fisheye lenses, which has been adopted in several works[12, 43]. The most widely used fisheye model [34] describes distortion as a function of the angular distance from the projection center:
where , and the distances of projected points to the image plane are normalized to 1. The 3D Gaussian Splatting method [38] assumes a standard perfect pinhole camera model and typically relies on COLMAP [59] to undistort images before reconstruction. To remove this constraint, recent methods [47, 51] adopt parametric models like Eq. 3 to extend 3D Gaussian Splatting techniques to fisheye images. However, these methods still depend heavily on camera calibration for accurate estimation and fix the projection in rasterization, limiting their generalizability to various camera types. Some works [51] introduce ray tracing into the rasterization pipeline and approximate Gaussian bounding using an icosahedron, which can potentially compromise rasterization efficiency. Other approaches [46, 4, 33] explore reconstruction from omnidirectional 360° panoramas, but the key difference is that panoramas require calibration to stitch two fisheye images together, which does not preserve raw geometric consistency at the stitching boundary. We address these limitations by introducing a hybrid distortion field that is compatible with the 3D Gaussian Splatting pipeline. Our experiments demonstrate that existing methods, such as Fisheye-GS [47] and ADOP [57], which incorporate traditional camera distortion models from Eqs. 2 and 3 into the rasterization process, are not expressive enough to handle the severe distortions present in large-FOV cameras.
@@
The bundle adjustment process can be extended to optimize camera lens parameters alongside poses, a process known as self-calibration [67, 56, 22, 31]. Camera calibration without a known calibration target is particularly challenging, as it relies on strong assumptions about scene structure and geometric priors to establish reliable correspondences [6, 14, 2]. Camera auto-calibration methods [67, 56, 22] extend this idea by deriving camera intrinsics from multi-view observations of unstructured scenes, an approach further advanced in recent studies [24, 23, 29, 21]. Several non-parametric models have been developed to ensure broad applicability across different camera and lens combinations [26, 13, 30, 45], while additional regularization is often required to maintain smooth underlying distortion [53]. With advances in differentiable rendering and rasterization pipelines [42, 64], recent works [57, 61, 36] have demonstrated that camera lens distortion can be optimized jointly with other parameters through a differentiable projection module. Prior works have also adapted NeRF for panoramic and fisheye-distorted inputs [28, 32, 62, 44]. These solutions typically rely on parametric models tailored for specific lenses, limiting their generalizability to a broader range of lens types. SCNeRF [36] models a residual projection matrix and residual raxel parameters [27], which are interpolated on a sub-sampled pixel grid. NeuroLens [61] optimizes lens parameters through an invertible neural network, while SC-OmniGS [33] optimizes camera parameters jointly with reconstruction but relies on calibrated captures. Building on insights from prior self-calibration methods, this work introduces a novel and efficient approach to modeling lens distortion, fully integrated with 3DGS [38].
@@ 3 Method
Given uncalibrated wide-angle captures, we aim to develop an algorithm that produces high-quality reconstructions using 3D Gaussians. Our method is designed to be robust against severe distortion in the peripheral regions of images and various wide-angle lens effects. First, we extend Gaussian Splatting to support a broader range of camera models, including fisheye lenses, as discussed in Sec. 3.2. To model lens distortion, we introduce a hybrid distortion field, enabling our approach to generalize across diverse real-world scenarios involving cameras with varying distortions. Second, we replace the traditional single-plane projection in 3DGS with a cubemap representation and introduce a distance-sorting strategy accordingly, as detailed in Sec. 3.3. Finally, we derive an efficient optimization for camera parameters, supporting self-calibration with distortion, as described in Sec. 3.4.
@@ 3.1 Background of Gaussian Splatting
A 3D Gaussian is parameterized by center , covariance , opacity , and color , represented using spherical harmonics. The 3D Gaussians are projected [68] into and , where is the affine approximation of the projection at point , parameterized by . For a pixel location , the RGB color is produced with alpha blending [42, 64]:
where is an ordered index of Gaussians sorted by depth, and depends on , , and . Finally, this set of 3D Gaussians is optimized using the L1 loss and D-SSIM [5] loss, along with adaptive control [38].
@@ 3.2 Lens Distortion Modeling
In this section, we extend the Gaussian Splatting technique to accommodate a broader class of camera lenses, including fisheye and wide-angle cameras, by modeling lens distortion. Lens distortion is typically captured by a distortion function defined in camera coordinates. A distortion function parameterized by maps pixel locations from a rectified image to locations in a distorted image. Ideally, the mapping should be: 1) expressive enough to model various lens distortions, 2) well-regularized so that it can be optimized together with the 3D scene, and 3) efficient, ensuring that it does not add significant computational overhead. While existing methods have explored using parametric camera models, grid-based methods, and deep-learning methods, none of these approaches perfectly satisfy all three criteria.
@@
where the optimizable parameters define an grid storing 2D vectors representing the distortion. The bilinear interpolation function is given by , where represents the bilinear interpolation weights at location x. Such a grid-based method is both expressive and efficient, as is sparse, and increasing the grid resolution allows for modeling more complex functions. However, the grid-based method lacks the smoothness required to model lens distortion properly, leading to overfitting and suboptimal solutions (Fig. 3 Top).
@@
An alternative way to model distortion is by using a neural network with an appropriate inductive bias. NeuroLens [61] proposes using an invertible ResNet [8] to represent non-linear lens distortions as a diffeomorphism. Specifically, the deformation mapping is modeled by a residual network:
where is a neural block parameterized by with a Lipschitz constant bounded by 1 (i.e., for all , , and ). represents a residual block with four linear layers. denotes the total number of blocks, which is 5 in our case, and the circle denotes function composition. Such constraints make the network invertible, and its inverse can be obtained using a fixed-point algorithm [8]. In the supplementary, we also provide additional comparisons and insights between iResNet and a regular ResNet.
While an invertible ResNet offers both expressiveness (i.e., the ability to model various lenses) and regularization, it is computationally prohibitive to apply it directly to 3DGS. To backpropagate gradients to the alpha-blending weights when rendering an image in Eq. 4, the computational graph for the backward passes of must be maintained for each Gaussian. This is infeasible due to the large number of 3D Gaussians in a single scene, often reaching millions and leading to out-of-memory errors. This limitation motivates us to develop a more efficient solution that leverages the inherent inductive bias of the invertible ResNet while maintaining computational efficiency.
@@
Given that the grid-based method is efficient yet tends to overfit, while the invertible ResNet has an appropriate inductive bias but is not efficient, we propose a hybrid method that combines the advantages of both. Specifically, we use the invertible ResNet to predict the flow field on a sparse grid and apply bilinear interpolation to each projected 2D Gaussian:
where is a sparse grid of fixed control points (pixel locations, where represents the resolution of control points rather than image resolution), and is an invertible ResNet parameterized by .
Unlike existing hybrid neural fields [52], where networks are applied after grid interpolation, our approach uses iResNet to predict displacement vectors on a sparse grid, with bilinear interpolation applied to produce a continuous displacement field, as shown at the bottom of Fig. 3. The advantage of this architecture is that we only need to compute the expensive forward and backward ResNet mappings for locations at , which scales with the grid resolution and is independent of the number of Gaussians in the scene. The additional operation required for each Gaussian is , which is computationally affordable and parallelizable.